A student dissolves calcium carbonate solids, CaCO
Answer #1

A) CaCO3 (s) <---> Ca2+ (aq) + CO32-(aq)

Now, the molar ratio of ions is 1:1, so we can say that the concentrations of both the ions is equal.

Now, Ksp = [Ca2+][CO32-].

Ksp = x . x

Ksp = X2

(Ksp)1/2 = X

X = (5.0 X 10-9)1/2 = 0.0000707106 = 7.07 x 10-5 moles/L

converting moles to g

moles x molecular weight = given weight

7.07x 10-5 mol X 100.0869 g/mol = 707.61x10-5 = 7.07 x 10-3 g/l

B) answer to this would take care of the common ion effect on molar solubility.

0.01mol/L of sodium carbonate

now,

[Ca2+] = x

[CO32-] = 0.01+x

now, solubility product equation is

Ksp = (x )(0.01+x)

5.0X 10-9 = 0.01x + X2

0.01x + X2 - 5.0X 10-9 =0

X2 +0.01x - 5.0X 10-9 =0

Solving quadratic

x = -0.01+/-[(0.01)2 - 4(1)(- 5.0X 10-9)]1/2/ 2

x = 4.99 x 10-7 mol/L

so, now the solubility is much lesser and hence the solid would precipitate out more quickly on addition of a common as seen here

C) CaO(s) + H2O(l) + Na2CO3(s) ---> CaCO3(s) + 2 NaOH(aq)

moles of CaO =0.0140/56 = 0.00025 moles

moles of Na2CO3 =0.0265/106 =0.00025 moles

moles of CaCo3 formed would be = 0.00025 moles

if ion product < Ksp no precipitation occurs

ion product is 6.25 x 10-8 and Ksp = 5.00 X 10-9

so, here precipitation will take place.